# When Context Is Not Enough: Modeling Unexplained Variability in Car-Following Behavior

### v1

@@ 4.2.2 Scenario-specific analysis and interpretability
    A reactive lengthscale that contracts during abrupt transitions and expands during smooth driving phases;
@@ 4.3 Quantitative Evaluations of Simulation Performance
RMSE quantifies the average deviation between the simulated and ground-truth trajectories for acceleration, speed, and gap distance, thus serving as a standard measure of prediction accuracy (Zhou et al.,, 2025). Formally, it is computed as:
@@
CRPS is employed to evaluate the quality of uncertainty quantification in one-dimensional stochastic forecasts. Unlike RMSE, which considers only the predicted mean, CRPS assesses the entire forecast distribution $F(x)$ against the observed outcome $x_{t}$, measuring both calibration and sharpness (Matheson and Winkler,, 1976):
[…]
@@
[…]
@@ 4.3.1 Impact of kernel design on simulation performance
[…]
@@
Table 2 summarizes the RMSE, CRPS, and Energy Score (ES) across acceleration, speed, and spacing for 10‐second stochastic simulations under four covariance structures: Gibbs kernel (nonstationary GP), Matérn(5/2) kernel (stationary GP), SE kernel (stationary GP), and White Noise kernel (i.i.d.). These results offer a comprehensive comparison of how kernel design influences both predictive accuracy and uncertainty quantification in car-following simulations.
@@
The White Noise model serves as a naive baseline, modeling each time step as conditionally independent. As expected, it yields the highest RMSE, CRPS, and ES values across all states, demonstrating that failure to model temporal correlations leads to degraded performance, especially in sequential tasks such as trajectory forecasting.
Figure 4: Error distributions under the DeepAR + Gibbs (nonstationary GP) configuration over all test simulations. Each row corresponds to a different metric. Top: RMSE, middle: CRPS, bottom: ES. And each column corresponds to a different state (acceleration, speed, spacing). Histograms are normalized to percentage frequency; dashed vertical lines denote the mean values reported in Table 2. Note the long tails in the spacing histograms, indicating that although the average gap error is low, extreme underestimations or overestimations can still occur.
@@ 4.3.2 Comparative model analysis and temporal correlation effects
Across all evaluation metrics, DeepAR combined with the nonstationary Gibbs kernel exhibits superior performance. Notably, it achieves the lowest RMSE across the three covariates, indicating its effectiveness in capturing abrupt and context-dependent changes in driver behavior. In terms of uncertainty quantification, the CRPS values are also consistently lower than those produced by models using stationary kernels, validating the Gibbs kernel’s capacity to adaptively characterize residual variability. The ES further confirms this advantage, all of which the values substantially outperform the corresponding metrics from models employing Matérn or SE kernels.
@@
5 Conclusion and Discussions
@@
Overall, this study provides strong evidence that integrating structured, nonstationary residual modeling into deep neural networks significantly improves simulation fidelity and interpretability. Beyond DeepAR, the modular nature of our approach, which is embodied in the general formulations of Equation 2 and Equation 3, makes it readily extensible to other neural architectures or kernel choices. This opens new avenues for probabilistic modeling of more general driving behaviors, particularly under heterogeneous or safety-critical conditions.

### v2

@@ 4.2.2 Scenario-specific analysis and interpretability
    A reactive lengthscale that contracts during rapid regime transitions and expands during smooth driving phases;
@@ 4.3 Quantitative Evaluations of Simulation Performance
RMSE quantifies the average deviation between the simulated and ground-truth trajectories for acceleration, speed, and gap distance, thus serving as a standard measure of prediction accuracy (Zhou et al., 2025). Formally, it is computed as:
@@
CRPS is employed to evaluate the quality of uncertainty quantification in one-dimensional stochastic forecasts. Unlike RMSE, which considers only the predicted mean, CRPS assesses the entire forecast distribution $F(x)$ against the observed outcome $x_{t}$, measuring both calibration and sharpness (Matheson and Winkler, 1976):
$\mathrm{CRPS}(x_{t}):=\int_{-\infty}^{+\infty}\left(F(x)-\mathbbm{1}\{x>x_{t}\}\right)^{2}dx,$ (25)
@@
where $\mathcal{P}$ is the predictive distribution, $\bm{x}$ is the observed vector of target variables, $\bm{X}^{(m)}$ are samples drawn from $\mathcal{P}$, $M$ is the number of samples, and $\lVert\cdot\rVert$ denotes the Euclidean norm. The ES penalizes discrepancies between the simulated samples and the ground truth while also penalizing inconsistencies within the simulated samples. This dual focus makes it an effective metric for evaluating joint predictive performance in multi-step stochastic simulations. Lower ES indicates better agreement between the predicted and observed multivariate distributions, reflecting the model’s ability to generate realistic and reliable forecasts.
@@ 4.3.1 Impact of kernel design on simulation performance
Metrics $\backslash$ Model DeepAR MLP (with delay embedding) Bayesian IDM (pooled)
@@
Table 2 summarizes the RMSE, CRPS, and Energy Score (ES) across acceleration, speed, and spacing for 10‐second stochastic simulations under four covariance structures: Gibbs kernel (nonstationary GP), Matérn(5/2) kernel (stationary GP), SE kernel (stationary GP), and White Noise kernel (i.i.d.). Compared to the simple white-noise baseline, the nonstationary Gibbs residual yields large absolute gains, while adding only a short-horizon covariance structure in simulation. These results offer a comprehensive comparison of how kernel design influences both predictive accuracy and uncertainty quantification in car-following simulations.
@@
The White Noise model serves as a baseline representing the standard independent error assumption commonly found in traditional calibration studies (Hoogendoorn and Hoogendoorn, 2010; Punzo et al., 2021). By modeling each time step as conditionally independent (i.e., assuming i.i.d. residuals), this configuration reflects the performance of models that neglect structured temporal dependencies. As expected, it yields the highest RMSE, CRPS, and ES values across all states, demonstrating that standard calibration practices may lead to degraded performance in sequential tasks such as trajectory forecasting due to the failure to model temporal correlations.
Figure 5: Error distributions under the DeepAR + Gibbs kernel (nonstationary GP) configuration over all test simulations. Each row corresponds to a different metric. Top: RMSE, middle: CRPS, bottom: ES. Each column corresponds to a different state (acceleration, speed, spacing). Histograms are normalized to percentage frequency; dashed vertical lines denote the mean values reported in Table 2. Note the long tails in the spacing histograms, indicating that although the average gap error is low, extreme underestimations or overestimations can still occur.
@@ 4.3.2 Comparative model analysis and temporal correlation effects
Across all evaluation metrics, DeepAR combined with the nonstationary Gibbs kernel exhibits superior performance. Notably, it achieves the lowest RMSE across the three covariates, indicating its effectiveness in capturing abrupt and context-dependent changes in driver behavior. In terms of uncertainty quantification, the CRPS values are also consistently lower than those produced by models using stationary kernels, validating the Gibbs kernel’s capacity to adaptively characterize residual variability. The ES further confirms this advantage, all values substantially outperform the corresponding metrics from models employing Matérn or SE kernels.
@@
4.4 Discussion: Framework Synthesis, Practical Implications, and Transferability
Framework synthesis.
This paper argues that a substantial portion of car-following variability remains unexplained even after conditioning on commonly used contextual variables, and that this residual exhibits temporally structured dependence that can change across driving regimes. Consistent with this view, replacing an i.i.d. disturbance with a temporally correlated residual improves both predictive accuracy and probabilistic consistency, with the largest gains when correlation structure is allowed to vary over time. These findings support a coupled mean–residual decomposition: a mean dynamics captures predictable interaction structure, while a stochastic residual process captures regime-dependent, temporally correlated deviations beyond the mean. The benefit arises from their synergy: improving the mean reduces systematic bias, whereas modeling residual dependence improves uncertainty representation and temporal realism.
Model-agnostic residual module and kernel choice.
Although we instantiate the mean dynamics using DeepAR, the nonstationary GP is defined on residuals and can be combined with a broad family of car-following models, including more behaviorally structured formulations (e.g., safety-constrained, kinematic, traffic-flow-inspired, and psycho-physical models). The kernel choice should reflect residual characteristics: a stationary SE kernel can suffice when residuals are smooth and stable within an episode; Matérn kernels are preferable for rougher or faster-varying residuals; and the nonstationary Gibbs kernel is most useful when correlation lengthscales and variance shift across regimes (e.g., transitions from free-flow to close following), where a single global lengthscale is inadequate.
Transferability under domain shift.
Because the framework learns both mean dynamics and context-dependent residual structure from data, performance depends on how well deployment conditions are represented by the training distribution. Under substantial domain shift (e.g., different road geometry, traffic rules, or driver populations), generalization is not guaranteed. A useful separation is what may transfer (shared regimes and stable physical meaning of context variables such as headway and relative speed) versus what may require adaptation (qualitatively different interaction patterns). In practice, lightweight adaptation can be implemented by fine-tuning a subset of parameters using a small target-domain calibration set, for example updating kernel modulation and output scales, or adapting the mean while retaining kernel structure when temporal dependence appears stable.
Measurement noise and evaluation.
Acceleration is often obtained by differentiating discrete trajectories, which can amplify measurement noise and distort residual autocorrelation if not accounted for. Our formulation partially addresses this by separating observation noise from the temporally correlated residual: the additive noise term absorbs unstructured measurement error, while the GP residual represents structured, lower-frequency deviations beyond the mean. In practice, preprocessing (e.g., smoothing or downsampling prior to differentiation) and plausibility checks (e.g., jerk statistics) are important; a principled extension is to treat acceleration as latent and introduce an explicit measurement model. Finally, evaluation should match downstream use: probabilistic scores such as CRPS complement point metrics for forecasting, while simulation requires additional checks of physical plausibility, stability, and reproduction of emergent stop-and-go phenomena.
Identifiability and noise decomposition.
Following the above discussion, a critical aspect of the proposed framework is the decomposition of the residual into a temporally correlated behavioral process $\delta(t)$ and an i.i.d. observation noise $\epsilon$. This separation is essential for model identifiability; mathematically, if both components were permitted to exhibit temporal correlation, the likelihood optimization would face a significant signal-separation challenge, struggling to distinguish between structured driver behavior and high-frequency measurement artifacts. By treating $\epsilon$ as a “nugget effect,” we absorb irreducible sensor noise and differentiation errors into the white-noise term. This formulation forces the GP kernel to prioritize the capture of lower-frequency, persistent behavioral dependencies rather than measurement artifacts. The same identifiability issue of the Bayesian method within this framework was previously discussed in Zhang and Sun (2024) and Zhang et al. (2024b).
Practical deployment and evaluation in microsimulation.
While our experiments validate short-horizon stochastic rollouts on naturalistic trajectories, an important next step is closed-loop microsimulation evaluation, where the model is used as the driving policy for multiple interacting vehicles and assessed via macroscopic outcomes (e.g., wave propagation, stop-and-go formation, and safety-related events). Moreover, broader cross-dataset validation is needed to quantify robustness under heterogeneous road types, regions, and driver populations. For simulator integration, a practical route is to use the selected mean model as the nominal car-following rule and inject a temporally correlated residual in a receding-horizon manner (apply the first-step acceleration each tick), enabling plug-in deployment without changing the simulator’s underlying vehicle dynamics.
Extensions to delayed, non-smooth, and multi-lane settings.
The residual-based construction provides a direct path to richer settings. For delayed-response models (e.g., GM-type), the mean can incorporate delayed covariates or an explicit reaction-time parameter while retaining the GP to capture remaining temporally correlated variability. For action-point or piecewise models with non-smooth accelerations, e.g., Wiedemann-type (Wiedemann, 1974), the mean can encode discontinuous updates, and the residual module can accommodate rougher deviations via Matérn or regime-dependent nonstationarity. Multi-lane extensions can be achieved by augmenting context with adjacent-lane states and maneuver cues (e.g., neighboring gaps, closing rates, lane-change indicators), allowing uncertainty to adapt during lane-changing and merging interactions.
5 Conclusion
@@
Overall, this study provides strong evidence that integrating structured, nonstationary residual modeling improves both simulation fidelity and uncertainty calibration for car-following. These capabilities support practical applications such as uncertainty-aware integration into microscopic traffic simulators and safety-critical connected and automated vehicles (CAV) analysis through risk-aware scenario evaluation and robustness testing under human-driving uncertainty.